OpenStudy (anonymous):

Simplify (4.3 x 105) (2.7 x 107)and write in scientific notation

OpenStudy (anonymous):

please answer my question

OpenStudy (anonymous):

i will give you a metal

OpenStudy (anonymous):

(4.3 x 105) (2.7 x 107) (451.5) (288.9) = 130438.35 In scientific notation = 1.3043835 x 10^5

OpenStudy (anonymous):

sorry It took so long I was busy :) @alison124

OpenStudy (anonymous):

\[ 1.3043835 \times10 ^{5}\]

imqwerty (imqwerty):

sry there was a netwrk error :) open brackets nd get-\[4.3\times 10^5 \times 2.7 \times 10^7 \] simplify 4.3 x 2.7=11.61 and 10^5 x 10^7=10^12 so we get- \[11.61 \times 10^\left( 12 \right)\]to write it in scientific notation we put decimal after the 1st place so it wuld be like-\[1.161\times 10 \times 10^\left( 12 \right) =>1.161 \times 10^\left( 13 \right)\]